we drive a distance of 1km at 16km/hr. Then we drive am additional distance of 1 km at 32 km/hr. what ia our average speed?
1
Expert's answer
2020-10-25T18:31:22-0400

We can find the average speed from the formula:


"v_{avg}=\dfrac{Total\ distance}{Total\ time\ elapsed}."

The total distance traveled is the sum of two distances: the initial distance "s_1" and the additional distance "s_2". Therefore, we can write:


"Total\ distance=s_1+s_2=1\ km+1\ km=2\ km."

The total time elapsed is the sum of two times: the time "t_1"that the car takes to travel the distance "s_1"and the time "t_2" that the car takes to travel the additional distance "s_2".

Therefore, we can write:

"Total\ time\ elapsed=t_1+t_2=\dfrac{s_1}{v_1}+\dfrac{s_2}{v_2},""Total\ time\ elapsed=\dfrac{1\ km}{16\ \dfrac{km}{hr}}+\dfrac{1\ km}{32\ \dfrac{km}{hr}}=0.094\ hr."

Finally, we can find the average speed:


"v_{avg}=\dfrac{Total\ distance}{Total\ time\ elapsed}=\dfrac{2\ km}{0.094\ hr}=21.3\ \dfrac{km}{hr}."

Answer:

"v_{avg}=21.3\ \dfrac{km}{hr}."